OpenStudy (anonymous):

Simplify each expression: 1- (-18) + (-11) 9 + 2 (16/4 - 5) -28/4+3 Show all work necessary to solve.

OpenStudy (anonymous):

1- (-18) + (-11) = 1 + 18 - 11 = 8 9 + 2 (16/4 - 5) = 9 + 2(4 - 5) = 9 + 2(-1) = 9 - 2 = 7 -28/4+3 = -7 + 3 = -4